OpenStudy (anonymous):

Can someone help me solve this problem 10^2x - 2 = 10^x

OpenStudy (saifoo.khan):

quadratic eq. u want to factor it or want the values of x?

OpenStudy (anonymous):

value of x

OpenStudy (anonymous):

\[10^{2x}-10^x-2=0\] as saifoo said a quadratic in \[10^x\]

OpenStudy (saifoo.khan):

x = 1.17 x = -0.17

OpenStudy (anonymous):

can you explain saifoo

OpenStudy (saifoo.khan):

u should solve it with the quadratic formula. As u r going to insert the values, u will automatically arrive to this answer.

OpenStudy (anonymous):

solve this as \[z^2-z-2=0\]

OpenStudy (anonymous):

where \[z=10^x\]

OpenStudy (anonymous):

get \[(z-2)(z+1)=0\] \[z=2\] \[z=-1\]

OpenStudy (anonymous):

so you have \[10^x=2\] \[x=\log(2)\] or \[10^x=-1\] which has no solution

OpenStudy (anonymous):

ok, thank you

OpenStudy (anonymous):

not sure what \[\log(2)\] is but you can find it with a calculator. you do not need the quadratic formula for this because it factors

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

\[10^{2x} - 2 = 10^{x} <=> (10^{x})^2 - 10^x - 2 = 0\]\[z = 1o^x\] so \[z^2 - z - 2 = 0\] the solutions are z = 4 and z = -2 \[10^x = 4 \]and \[10^x = -2\]the second equation has no solution 10^x = 4 => x=log4